OpenStudy (anonymous):

Match each equation with the scenario it best represents. Match Term Definition x minus three equals seven A) Reina completed 7 homework questions. She still has ten questions to complete. seven plus x equals ten B) At the party, 7 people arrived before the party began. By the end of the party, there were ten people there. x plus three equals seven C) Audra ate 3 slices of bread and now has 7 left. x minus seven equals ten D) Jackson's hair grew 3 inches last month and is now 7 inches long.

OpenStudy (mathmale):

I've spent several minutes looking over your posted problem and still find it a bit confusing, not because of the algebra involved, but because it's just plain hard to read. I think you probably were given this problem with the info arranged in columns for easy reading. I'd suggest you type the four algebraic expressions on the left and use the space bar to create a new column on the right: 1) x minus three equals seven A) Reina completed 7 homework questions. She still has ten questions to complete. 2) seven plus x equals ten B) It would be a lot easier for me, and perhaps for you, if you'd arrange the info neatly so that it's easy to match up the algebra with the four statements. I know this must sound picky to you, but would you please do it? All you have to do is to copy the problem as you originally typed it in and then work on the format using the space bar and Enter key.

OpenStudy (anonymous):

Term Definition x - 3 = 7 A ) Reina completed 7 homework questions. She still has ten questions to complete

OpenStudy (anonymous):

Term 7 + x = 10 Definition At the party, 7 people arrived before the party began. By the end of the party, there were ten people there

OpenStudy (anonymous):

the first one is a the second one is b

OpenStudy (mathmale):

Sorry that it's a bit hard to format anything using only the space bar and enter key. Let's take an alternative approach. Look at x minus 3 equals 7. Then look through your four choices and choose the one that you think best fits this statement, and why.

OpenStudy (mathmale):

Hint: Look at (C). We're not told how many slices of bread this girl started with. After she eats 3 slices, she has 7 slices left. That's x-3=7. Match up the first algebraic expression with (c). Now you have only 3 more expressions and sentences to match up.